OpenStudy (anonymous):

factor 6x^2 + 5x - 1 ?

OpenStudy (anonymous):

-1 and 1/6

OpenStudy (ajprincess):

Find two factors of -6 whose sum is 5. can u?

OpenStudy (anonymous):

i used the quadratic formula

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

mmhm no problem

OpenStudy (ajprincess):

you are nt asked to find the roots but to factorise @Andres07

OpenStudy (anonymous):

um, you can factor a quadratic like that?

OpenStudy (anonymous):

http://www.purplemath.com/modules/solvquad.htm

OpenStudy (ajprincess):

bt the factors of a quadratic are given in this form (x-a)(x-b)

OpenStudy (anonymous):

i could have solved it by completing the square. i could have solved it many ways. by finding coefficients of 6 would make your life a lot harder

OpenStudy (ajprincess):

6x^2+5x-1 =6x^2+6x-x-1 =6x(x+1-(x+1) =(6x-1)(x+1)

OpenStudy (anonymous):

Okay, has your teacher ever let you leave an answer like that? You're supposed to solve for x. Always

OpenStudy (anonymous):

@emkaye16 do you have to solve for x?

OpenStudy (ajprincess):

that's only when u r given a question like this solve 6x^2+5x-1=0

OpenStudy (ajprincess):

@andres07.

OpenStudy (anonymous):

you never offended me.

OpenStudy (anonymous):

at the moment no. If the question asks to solve for x, you solve for x. to factor, you leave the form x = -1 and 1/6 as (x + 1)(6x - 1). Both of you are right.